OpenStudy (anonymous):

Which are the solutions to the equation |2x 14| 12 = 12? (Points : 4) x = 10 or x = 7 x = 5 or x = 7 x = 5 or x = 2 x = 7 only

OpenStudy (jdoe0001):

|2x 14| 12 = 12? <--- is there like a + or a - somewhere? is it |2x 14| +12 = 12? or |2x 14| - 12 = 12?

OpenStudy (anonymous):

|2x - 14| -12 = 12?

OpenStudy (jdoe0001):

\(\bf |2x - 14| -12 = 12 \implies |2x - 14| = 12+12\\ \quad \\ \quad \\ |2x - 14| = 24 \implies \begin{cases} +1(2x - 14) = 24\\\quad \\ \bf -1(2x - 14) = 24 \end{cases}\)

OpenStudy (jdoe0001):

so, you'd get 2 answers, 2 values for "x", one from each scenario/case/function

OpenStudy (anonymous):

thanks i got it